OpenStudy (anonymous):

What is the solution set of x2 + 5x - 5 = 0

OpenStudy (anonymous):

[-5+/- sqrt (25+20)]/2 -2.5+/- 3sqrt(5)

OpenStudy (aroub):

can u do it by factoring too ( )( ) ????

OpenStudy (dumbcow):

no factoring wont work -5+1 = -4 -1 + 5 = 4